# QSpec: Speculative Decoding with Complementary Quantization Schemes

### v1

@@ Abstract
Quantization has been substantially adopted to accelerate inference and reduce memory consumption of large language models (LLMs). While activation-weight joint quantization speeds up the inference process through low-precision kernels, we demonstrate that it suffers severe performance degradation on multi-step reasoning tasks, rendering it ineffective. We propose a novel quantization paradigm called QSpec, which seamlessly integrates two complementary quantization schemes for speculative decoding. Leveraging nearly cost-free execution switching, QSpec drafts tokens with low-precision, fast activation-weight quantization, and verifies them with high-precision weight-only quantization, effectively combines the strengths of both quantization schemes. Compared to high-precision quantization methods, QSpec empirically boosts token generation throughput by up to 1.80×1.80$\times$1.80 × without any quality compromise, distinguishing it from other low-precision quantization approaches. This enhancement is also consistent across various serving tasks, model sizes, quantization methods, and batch sizes. Unlike existing speculative decoding techniques, our approach reuses weights and the KV cache, avoiding additional memory overhead. Furthermore, QSpec offers a plug-and-play advantage without requiring any training. We believe that QSpec demonstrates unique strengths for future deployment of high-fidelity quantization schemes, particularly in memory-constrained scenarios (e.g., edge devices).
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities across various domains, including mathematics, coding, and planning (Shao et al., 2024b; Guo et al., 2024a; Huang et al., 2024). Nonetheless, their immense scales pose substantial challenges for deployment due to high memory and computational demands, especially in resource-limited scenarios (e.g., inference on edge devices). Quantization has been an effective compression technique to facilitate LLM inference with limited resources (Lin et al., 2024a; Ashkboos et al., 2024; Zhao et al., 2024c; Lin et al., 2024b). By converting high-precision values (e.g., FP16) into their lower-precision counterparts (e.g., INT4), quantization effectively lowers memory and computational requirements, allowing for larger serving batches and model sizes. Furthermore, the reduced memory footprint boosts token generation throughput by accelerating the typically memory-bound autoregressive decoding process (Zhao et al., 2024a).
Based on the quantized objects, recent quantization algorithms can be broadly classified into two categories: weight-only and WXAX: (1) Weight-only quantization, represented by W4A16 (Lin et al., 2024a), quantizes model weights to low precision (e.g., 4-bit) for storage, and then dequantizes them to a higher precision (i.e., FP16) during inference; (2) WXAX methods, such as W4A4 (Ashkboos et al., 2024; Zhao et al., 2024c) and W8A8 (Xiao et al., 2023), simultaneously quantize both weights and activations, and leverage low-precision hardware support for faster execution without dequantizing them to higher precision. Nevertheless, WXAX schemes generally suffer model performance degradation due to more low-precision activations used (as verified in Sec. 2). This poses a tough trade-off between efficacy and efficiency, raising the question:
@@
Considering the comparable performance claims on recent W4A4 methods (Zhao et al., 2024c; Ashkboos et al., 2024), we first contend that their conclusions are biased due to limited evaluation tasks, and W4A4 still experiences significant performance drops when compared to their higher-precision activation counterparts. Specifically, while W4A4 schemes such as Atom (Zhao et al., 2024c) and QuaRot (Ashkboos et al., 2024) perform well on general tasks, such as PIQA (Bisk et al., 2020), Winogrande (Sakaguchi et al., 2019) and ARC Clark et al. (2018), they demonstrate notable performance declines in multi-step reasoning, particularly on mathematical and coding benchmarks (Xiong et al., 2024; Guo et al., 2024b) (shown in Table 1). This raises concerns about the comprehensiveness of evaluation and emphasizes the necessity of incorporating multi-step reasoning tasks into quantization assessment.
@@
We evaluate the generation quality and end-to-end serving throughput of QSpec against W4A4 and W4A16 schemes across multiple datasets, model sizes, quantization methods, and batch sizes. Empirically, QSpec preserves memory consumption and generation quality compared to W4A16, while offering a high acceptance rate and up to 1.80×1.80$\times$1.80 × higher token generation throughput, thereby mitigating the efficiency-efficacy trade-off of existing quantization methods. Notably, for multi-step reasoning tasks such as MATH (Hendrycks et al., 2021), QSpec fully compensates for an up to 51.11% decline in generation quality observed with existing W4A4 methods. Furthermore, QSpec provides plug-and-play compatibility without any training requirements, and can be seamlessly applied to any existing models, delivering superior performance with minimal effort.
@@
  3. We propose QSpec, synergizing two complementary weight-shared quantization schemes with speculative decoding, alleviating the efficiency-efficacy trade-off of quantization.
  4. Our empirical results reveal up to 1.80×1.80$\times$1.80 × acceleration without any quality sacrifice across diverse settings. Alongside consistent memory usage, these advantages promise QSpec for high-fidelity quantization deployment, especially in memory-constrained scenarios.
@@ 2.1 Compromised Performance of Activation Quantization
Table 1: Performance of Atom-based quantization schemes with different weight and activation precision across diverse tasks. “Acc”, “PPL” and “EM” stand for accuracy, perplexity, and exact match, respectively, with arrows indicating their positive trends.
@@
State-of-the-art (SOTA) activation-weight joint quantization methods, like Atom (Zhao et al., 2024c) and QuaRot (Ashkboos et al., 2024), achieve notable speed-ups with negligible performance loss compared to weight-only ones. However, we argue that this conclusion is skewed by the limited evaluation benchmarks, which fail to capture the negative impacts of activation quantization.
To substantiate this claim, we conduct experiments on Llama-3-8B-Instruct models (Dubey et al., 2024) quantized with W16A16, W4A16, and W4A4 methods across four task datasets: PIQA (Bisk et al., 2020), WikiText-2 (Merity et al., 2016), GSM8K (Cobbe et al., 2021), and MBPP (Austin et al., 2021). PIQA is a two-choice commonsense reasoning benchmark for physical knowledge, evaluated using classification accuracy, while WikiText-2 comprises a collection of high-quality Wikipedia articles, assessed for language fluency via perplexity (Jelinek et al., 1977). Both are commonly adopted in current quantization evaluations. GSM8K includes diverse grade school mathematical problems, evaluated by “exact match” metrics; MBPP focuses on crowd-sourced Python programming challenges, assessed by accuracy. Unlike the former two benchmarks, both GSM8K and MBPP necessitate auto-regressive multi-step reasoning abilities. While these critical abilities are propelled by the rapid advancement in LLMs recently, they have not yet been integrated into mainstream quantization evaluation. As shown in Table 1, Atom-based quantization schemes show comparable (and even superior sometimes) performance to W16A16 across commonly adopted tasks such as on PIQA and WikiText-2, aligning with the claims in Zhao et al. (2024c). However, W4A4 suffers a nearly 30% average performance decline on complex reasoning tasks (i.e., on MBPP and GSM8K), whereas W4A16 only experiences about 4%. This indicates that activation quantization leads to several times more performance degradation on multi-step reasoning tasks, despite the improved efficiency. Besides, the performance trend observed on multi-step reasoning tasks shows a stronger correlation with quantization precision than perplexity does, validating their adequacy in assessing quantization performance.
@@ 3 Method
Targeting an efficient quantization scheme without sacrificing performance or increasing memory consumption, we propose a new quantization paradigm called speculative decoding with complementary quantization execution (QSpec). As shown in Figure 1, QSpec employs and-verpipeline for next-token prediction with varying activation precisions and shared low-precision quantized weights, instead of a constantsingleaquantization scheme. The details are elaborated below, adhering to the core component breakdown of regular speculative decoding (Leviathan et al., 2023).
@@ 3.1 QSpec
To maintain high reproducibility, both low-precision and high-precision activation quantization schemes utilize greedy decoding throughout the generation process. This means that one drafted tokens $\hat{t}_{i+j}$ is accepted as ${t}_{i+j}$ only when the top-1 tokens from ${p}_{i+j}$ and $\hat{p}_{i+j}$ coincide; otherwise, this token is rejected. Nonetheless, we claim that alternative strategies, as outlined in Leviathan et al. (2023), can be directly applied to our method due to the similarities in the framework. Figure 3 illustrates a mini-sample of this cycle with the draft token length $\gamma=4$. The model initially speculates four tokens using W4A4 scheme. Subsequently, adhering to a predefined acceptance policy, it accepts all drafted tokens after verifying them through the W4A16 scheme. In the second loop, however, only the first two tokens are accepted. A new token “is” is directly derived from the prediction probability of W4A16 scheme, and another draft-verify cycle will commence from the ninth token.
@@
To further reduce memory consumption, QSpec overwrites KV caches of low-precision activation quantization with those of high-precision method. Specifically, compared to W4A4, W4A16 is expected to yield a higher quality FP16 KV cache due to higher activation precision. Alongside the shared weights, this naturally allows overwriting the low-quality KV caches generated by W4A4 with those from W4A16 for accepted tokens after each validation phase. This enables W4A4 to condition on high-quality KV caches for subsequent autoregressive generation, and saves the memory occupation of W4A4 KV caches, despite requiring negligible buffer space for temporarily storing. To some extent, this operation aligns with the settings of attention kernels in prior works (Zhao et al., 2024c; Shao et al., 2024a; Ashkboos et al., 2024), where INT4 KV caches are typically dequantized to FP16 before or during precision-sensitive attention operations to ensure accurate computations.
3.2 Advantage Analysis
@@
As shown in Table 2, we compare QSpec with individual quantization schemes (i.e., W4A4 and W4A16) as well as speculative decoding across the dimensions of memory, computation, and generation. QSpec offers several key advantages over these methods, detailed as follows:
  1. Memory-efficient. Quantization is often motivated by memory constraints, rendering regular speculative decoding unsuitable due to the additional memory allocation for the weights and KV caches of the draft model. However, QSpec addresses these memory overheads by sharing weights and overwriting KV caches, aligning with the costs associated with standalone high-precision activation quantization.
  2. No efficiency-efficacy trade-off. Leveraging the speculative decoding framework, QSpec achieves efficiency gains without any quality sacrifice, thereby avoiding the trade-off between efficiency and efficacy. In contrast, individual quantization methods either endure significant performance degradation or accept reduced inference speed.
  3. High acceptance rate. The shared weights inherently enable a strong similarity between the two quantization methods. Besides, the KV cache overwriting further enhances the consistency of subsequent predictions. Both factors collectively contribute to a high token acceptance rate of QSpec.
  4. Plug-and-play compatibility. Compared to individual quantization schemes, QSpec simply integrates an acceptance policy and a KV cache overwriting operation. This allows QSpec to be swiftly implemented based on existing quantization codes without extensive modifications. Furthermore, QSpec operates without additional training or classifiers, enabling its direct application to any existing model for enhanced inference efficiency.

### v2

# Juntao Zhao1,*    Wenhao Lu2,†    Sheng Wang1,†    Lingpeng Kong1    Chuan Wu1

@@ Abstract
Quantization has been substantially adopted to accelerate inference and reduce memory consumption of large language models (LLMs). While activation-weight joint quantization speeds up the inference process through low-precision kernels, we demonstrate that it suffers severe performance degradation on multi-step reasoning tasks, rendering it ineffective. We propose a novel quantization paradigm called QSpec, which seamlessly integrates two complementary quantization schemes for speculative decoding. Leveraging nearly cost-free execution switching, QSpec drafts tokens with low-precision, fast activation-weight quantization, and verifies them with high-precision weight-only quantization, effectively combines the strengths of both quantization schemes. Compared to high-precision quantization methods, QSpec empirically boosts token generation throughput by up to 1.64×1.64$\times$1.64 × without any quality compromise, distinguishing it from other low-precision quantization approaches. This enhancement is also consistent across various serving tasks, model sizes, quantization methods, and batch sizes. Compared to state-of-art speculative decoding methods, our approach reuses weights and the KV cache, avoiding extra memory overhead while achieving up to 1.55×1.$5$ speedup in batched serving with a high acceptance rate. Furthermore, QSpec offers a plug-and-play advantage without requiring any training. We believe that QSpec demonstrates unique strengths for future deployment of high-fidelity quantization schemes, particularly in memory-constrained scenarios (e.g., edge devices).
@@ 1   Introduction
Large language models (LLMs) have demonstrated remarkable abilities across various domains, including mathematics, coding, and planning (Shao et al., 2024b; Guo et al., 2024a; Huang et al., 2024). Nonetheless, their immense scales pose substantial challenges for deployment due to high memory and computational demands, especially in resource-limited scenarios (e.g., inference on edge devices). Quantization has been an effective compression technique to facilitate LLM inference with limited resources (Lin et al., 2024a; Ashkboos et al., 2024; Zhao et al., 2024b; Lin et al., 2024b). By converting high-precision values (e.g., FP16) into their lower-precision counterparts (e.g., INT4), quantization effectively lowers memory and computational requirements, allowing for larger serving batches and model sizes. Furthermore, the reduced memory footprint boosts token generation throughput by accelerating the typically memory-bound autoregressive decoding process (Zhao et al., 2024a).
Based on the quantized objects, recent quantization algorithms can be broadly classified into two categories: weight-only and WXAX: (1) Weight-only quantization, represented by W4A16 (Lin et al., 2024a), quantizes model weights to low precision (e.g., 4-bit) for storage, and then dequantizes them to a higher precision (i.e., FP16) during inference; (2) WXAX methods, such as W4A4 (Ashkboos et al., 2024; Zhao et al., 2024b) and W8A8 (Xiao et al., 2023), simultaneously quantize both weights and activations, and leverage low-precision hardware support for faster execution without dequantizing them to higher precision. Nevertheless, WXAX schemes generally suffer model performance degradation due to more low-precision activations used (as verified in Sec. 2). This poses a tough trade-off between efficacy and efficiency, raising the question:
@@
Considering the comparable performance claims on recent W4A4 methods (Zhao et al., 2024b; Ashkboos et al., 2024), we first contend that their conclusions are biased due to limited evaluation tasks, and W4A4 still experiences significant performance drops when compared to their higher-precision activation counterparts. Specifically, while W4A4 schemes such as Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024) perform well on general tasks, such as PIQA (Bisk et al., 2020), Winogrande (Sakaguchi et al., 2019) and ARC Clark et al. (2018), they demonstrate notable performance declines in multi-step reasoning, particularly on mathematical and coding benchmarks (Xiong et al., 2024; Guo et al., 2024b) (shown in Table 1). This raises concerns about the comprehensiveness of evaluation and emphasizes the necessity of incorporating multi-step reasoning tasks into quantization assessment.
@@
We evaluate the generation quality and end-to-end serving throughput of QSpec against W4A4 and W4A16 schemes across multiple datasets, model sizes, quantization methods, and batch sizes. Empirically, QSpec preserves memory consumption and generation quality compared to W4A16, while offering a high acceptance rate and up to 1.64×1.64$\times$1.64 × higher token generation throughput, thereby mitigating the efficiency-efficacy trade-off of existing quantization methods. Notably, for multi-step reasoning tasks such as MATH (Hendrycks et al., 2021), QSpec fully compensates for an up to 51.11% decline in generation quality observed with existing W4A4 methods. Furthermore, QSpec provides plug-and-play compatibility without any training requirements, and can be seamlessly applied to any existing models, delivering superior performance with minimal effort.
@@
  3. We propose QSpec, the first quantization paradigm that synergizing two complementary weight-shared quantization schemes with speculative decoding, alleviating the efficiency-efficacy trade-off of quantization. We demonstrate its superior efficiency in quantized batch-serving compared to state-of-the-art speculative decoding approaches theoretically and empirically.
  4. Our empirical results reveal up to 1.64×1.64$\times$1.64 × acceleration without any quality sacrifice across diverse settings. Alongside consistent memory usage, these advantages promise QSpec for high-fidelity quantization deployment, especially in memory-constrained scenarios.
@@ 2.1   Compromised Performance of Activation Quantization
Table 1: Performance of Atom-based quantization schemes with different weight and activation precision across diverse tasks. “Acc”, “PPL” and “EM” stand for accuracy, perplexity, and exact match, respectively, with arrows indicating their positive trends. “W16A16” refers to standard FP16 inference, where both weights and activations are represented in FP16 precision.
@@
State-of-the-art (SOTA) activation-weight joint quantization methods, like Atom (Zhao et al., 2024b) and QuaRot (Ashkboos et al., 2024), achieve notable speed-ups with negligible performance loss compared to weight-only ones. However, we argue that this conclusion is skewed by the limited evaluation benchmarks, which fail to capture the negative impacts of activation quantization.
To substantiate this claim, we conduct experiments on Llama-3-8B-Instruct models (Dubey et al., 2024) quantized with W16A16, W4A16, and W4A4 methods across four task datasets: PIQA (Bisk et al., 2020), WikiText-2 (Merity et al., 2016), GSM8K (Cobbe et al., 2021), and MBPP (Austin et al., 2021). PIQA is a two-choice commonsense reasoning benchmark for physical knowledge, evaluated using classification accuracy, while WikiText-2 comprises a collection of high-quality Wikipedia articles, assessed for language fluency via perplexity (Jelinek et al., 1977). Both are commonly adopted in current quantization evaluations. GSM8K includes diverse grade school mathematical problems, evaluated by “exact match” metrics; MBPP focuses on crowd-sourced Python programming challenges, assessed by accuracy. Unlike the former two benchmarks, both GSM8K and MBPP necessitate auto-regressive multi-step reasoning abilities. While these critical abilities are propelled by the rapid advancement in LLMs recently, they have not yet been widely integrated into mainstream quantization evaluation. As shown in Table 1, Atom-based quantization schemes show comparable performance to W16A16 across commonly adopted tasks such as on PIQA and WikiText-2, aligning with the claims in Zhao et al. (2024b). However, W4A4 suffers a nearly 30% average performance decline on complex reasoning tasks (i.e., on MBPP and GSM8K), whereas W4A16 only experiences about 4%. This indicates that activation quantization leads to several times more performance degradation on multi-step reasoning tasks, despite the improved efficiency. Besides, the performance trend observed on multi-step reasoning tasks shows a stronger correlation with quantization precision than perplexity does, validating their adequacy in quantization evaluation.
@@ 3   Method
Targeting an efficient quantization scheme without sacrificing performance or increasing memory consumption, we propose a new quantization paradigm called speculative decoding with complementary quantization execution (QSpec). As shown in Figure 1, QSpec employs a draft-verify pipeline for next-token prediction with varying activation precisions and shared low-precision quantized weights, instead of a single quantization scheme. The details are elaborated below, adhering to the core component breakdown of regular speculative decoding (Leviathan et al., 2023).
@@ 3.1   QSpec
To maintain high reproducibility, both low-precision and high-precision activation quantization schemes employ greedy decoding. This means that one drafted tokens $\hat{t}_{i+j}$ is accepted as ${t}_{i+j}$ only when the top-1 tokens from ${p}_{i+j}$ and $\hat{p}_{i+j}$ coincide; otherwise, this token is rejected. Nonetheless, we claim that alternative strategies, as outlined in Leviathan et al. (2023), can be directly applied to our method due to the similarities in the framework. Figure 3 illustrates a mini-sample of this cycle with the draft token length $\gamma=4$. The model initially speculates four tokens using W4A4 scheme. Subsequently, adhering to a predefined acceptance policy, it accepts all drafted tokens after verifying them through the W4A16 scheme. In the second loop, however, only the first two tokens are accepted. A new token “is” is directly derived from the prediction probability of W4A16 scheme, and another draft-verify cycle will commence from the ninth token.
@@
To further reduce memory consumption, QSpec overwrites KV caches of low-precision activation quantization with those of high-precision method. Specifically, compared to W4A4, W4A16 is expected to yield a higher quality FP16 KV cache due to higher activation precision. Alongside the shared weights, this naturally allows overwriting the low-quality KV caches generated by W4A4 with those from W4A16 for accepted tokens after each validation phase. This enables W4A4 to condition on high-quality KV caches for subsequent autoregressive generation, and saves the memory occupation of W4A4 KV caches, despite requiring negligible buffer space for temporarily storing. To some extent, this operation aligns with the settings of attention kernels in prior works (Zhao et al., 2024b; Shao et al., 2024a; Ashkboos et al., 2024), where INT4 KV caches are typically dequantized to FP16 before or during precision-sensitive attention operations to ensure accurate computations. Note that KV compression (Hooper et al., 2024; Kang et al., 2024) is fully orthogonal to our method, e.g., it is also possible to use 4-bit KV cache for W4A16.
Figure 4: Analysis of QSpec advantages: (a) Tree-structured drafting architecture, (b) Roofline model showing computational efficiency transitions (memory-bound $\rightarrow$ compute-bound) with batch size scaling, and (c) Per-valid-token latency composition comparing the change of extra computational overhead and accepted token numbers with scaling batch size. SD denotes tree-structured speculative decoding.
@@
As shown in Table 2, we compare QSpec against individual quantization configurations (i.e., W4A4 and W4A16) and speculative decoding in terms of memory, computation, and generation. QSpec provides four key benefits:
(1) Memory-Efficient. By sharing weights and overwriting KV caches, QSpec incurs memory costs on par with standalone high-precision activation quantization. (2) No Efficiency–Efficacy Trade-off. QSpec leverages speculative decoding to boost efficiency while preserving output quality, avoiding the compromises seen in conventional quantization. (3) Plug-and-Play Compatibility. QSpec requires only an acceptance policy and a cache-overwriting step, requiring no additional training or classifiers. Hence, it can be quickly integrated into existing quantization pipelines. (4) High Acceptance Rate. Common weights and KV cache overwriting ensure consistent predictions, leading to a high rate of token acceptance.
3.2   Advantages of High Acceptance Rate
A key advantage of QSpec lies in its superior acceptance rate. While state-of-the-art speculative decoding approaches such as EAGLE Li et al. (2024a) and Medusa Cai et al. (2024) rely on compressed draft models and tree-structured drafting to balance speed and accuracy, QSpec achieves high performance without such compromises. We analyze how QSpec demonstrates superior performance compared to these methods in quantized settings below.
Tree-Structured Drafting. As shown in Figure 4 (a), tree-structured drafting (Chen et al., 2024; Miao et al., 2024; Li et al., 2024a, b) is a widely adopted technique to improve acceptance rates in speculative decoding. Instead of generating a single draft token at a time using the draft model $\mathcal{M}_{d}$, $\mathcal{M}_{d}$ select $k$ (i.e., topk) tokens and infers $\gamma$ times to form a draft tree with depth $\gamma$. The target model $\mathcal{M}$ then verifies the tokens using masked tree attention, with the exact verification strategy depending on the sampling method (Cai et al., 2024; Miao et al., 2024). Under greedy sampling, the highest-probability token at each position is selected, forming the longest branch with the common prefix as the accepted sequence. In this tree, root node presents the past sequence, any token $t_{i}$ has a path from the root node $r$ to $t_{i}$, denoted by $Path(r,t_{i})$, consisting of ancestors $a_{1},\ldots,a_{j}$.
Expected average accepted token number. Let $p_{a}(t)$ be the probability of accepting a token $t$. The average accepted token number is given by:
$H(k):=\sum_{l}^{\gamma}\sum_{i=1}^{k^{l}}\prod_{j\in Path(r,t_{i})}p_{a}(a_{j}),$ (2)
where $k$ controls the tree’s branching factor (width) . Larger $k$ indicates larger expected token acceptance. When $k=1$, this method reduces to standard speculative decoding.
Cost Analysis. Denote by $C_{d}(\cdot)$ and $C_{p}(\cdot)$ the computation cost for draft and verify, The first argument indicates the number of sequences, and the second argument is the prefix sequence length. For a prefix with length $s$, we have per-valid-token cost $v$ by dividing the total cost by the number of accepted tokens:
$v&=\frac{\overbrace{C_{d}(1;s)+\cdots+C_{d}\bigl{(}k^{\gamma{}-1};s+\gamma{}-1\bigr{)}}^{\text{Drafting cost}}\;+\;\overbrace{C_{p}\bigl{(}k^{\gamma{}-1};s+\gamma{}\bigr{)}}^{\text{Verify cost}}}{\underbrace{H(k)}_{\text{Accepted tokens}}}\sim\frac{C(k)}{H(k)}.\end{aligned}$
Batch serving efficiency of QSpec. From Equation 3, efficient tree-structured drafting benefits from less extra computational overhead $\Delta C(k)_{\text{SD}}$ against acceptance rate gains $\Delta H(k)$, as illustrated in Figure 4(c). QSpec advances this paradigm in batched serving. As shown in Figure 4(b), QSpec employs W4A4 quantization for drafting. While initially exhibiting two limitations - suboptimal hardware utilization during the drafting and higher computational costs compared to small/pruned models - these challenges diminish with increased batch sizes. Beyond critical batch size $bs^{*}$, QSpec’s draft phase transits from memory-bound to compute-bound, achieving superior FLOP/s with low-precision. Also, by eschewing penalties for acceptance rate maintenance, QSpec reduces costs by a factor of $k$. As a result, in Figure 4(c), $\Delta C(k)_{\textsc{QSpec}{}}$ curve maintains a consistently shallower slope than $\Delta C(k)_{\text{SD}}$ for $bs>bs^{*}$, enabling higher token generation throughput. This observation applies to both static (e.g., SpecInfer (Miao et al., 2024), Sequoia (Chen et al., 2024)) and dynamic (Wang et al., 2024; Li et al., 2024b) tree-structured drafting methods, all of which exhibit similar behaviors.